Which equation could be used to solve the following?
A pizza is divided into 8 slices. Matt ate 3 of the 8 pizza slices. He paid $3.75 for his share of the cost. How much did the whole pizza cost?
A.
B.
C.
D.
* I got A
Which equation could be used to solve the following?
Find three consecutive integers such that 3 times the smallest is 15 more than the sum of the largest and next to largest.
A. 3x = 15 + x + x + 1 + x + 2
B. 3(x + 1) = 15 + x + x + 2
C. 3x = 15 + x + 1 + x + 2
D. 3x = 15 + x + x + 1
* I got D
Which equation could be used to solve the following?
The length of a rectangle is 3 more than twice the width. What are the dimensions if the perimeter is 36?
A. w + 2w + 3 = 36
B. 2w + 2w + 3 = 36
C. 2w + 2(w + 3) = 36
D. 2w + 2(2w + 3) = 36

When solving an equation, what is the ultimate goal?
A. to isolate the variable
B. to remove fractions
C. to remove negative signs
D. to get all numbers on the right side of the equal sign

Answer by Alan3354(69443)

Which equation could be used to solve the following?
A pizza is divided into 8 slices. Matt ate 3 of the 8 pizza slices. He paid $3.75 for his share of the cost. How much did the whole pizza cost?
A.
B.
C.
D.
* I got A
A looks good, tho it didn't paste very well.
---------------------------
Which equation could be used to solve the following?
Find three consecutive integers such that 3 times the smallest is 15 more than the sum of the largest and next to largest.
A. 3x = 15 + x + x + 1 + x + 2
B. 3(x + 1) = 15 + x + x + 2
C. 3x = 15 + x + 1 + x + 2
D. 3x = 15 + x + x + 1
* I got D
C is it, the largest and next to largest = (x+1) + (x+2). D adds the smallest and the middle number.
--------------------------
Which equation could be used to solve the following?
The length of a rectangle is 3 more than twice the width. What are the dimensions if the perimeter is 36?
A. w + 2w + 3 = 36
B. 2w + 2w + 3 = 36
C. 2w + 2(w + 3) = 36
D. 2w + 2(2w + 3) = 36
C is it. P = 2L + 2W = 2W + 2(W+3)
-----------------------------
When solving an equation, what is the ultimate goal?
A. to isolate the variable
B. to remove fractions
C. to remove negative signs
D. to get all numbers on the right side of the equal sign
A. Get the variable on one side of the equal sign by itself.